### v1

@@ 4 Equivalent characterizations of the Aubin property
It should be emphasized that, for (1.1) with ${\mathcal{K}}$ being an arbitrary $C^{2}$-cone reducible set, it is still unknown if the strong regularity of the KKT system (1.7) is equivalent to the constraint nondegeneracy (2.13) combined with a certain second-order optimality condition. The currently known cases include the nonlinear programming [12], the NLSOCP [4], the NLSDP problem [42] and a composite matrix programming regarding matrix eigenvectors [8].
@@ 5 Conclusions
In this paper, we prove that at a locally optimal solution to the nonlinear semidefinite programming problem (1.8), the Aubin property of ${\mathbb{S}}_{\rm KKT}$ (1.6) is equivalent to the strong second-order sufficient condition plus the constraint nondegeneracy. This enables us to derive a series of equivalent characterizations of the Aubin property, which includes the strong regularity of the Karush-Kuhn-Tucker system (1.3). As a byproduct, for nonlinear semidefinite programming this paper answers the open question posed in [10, Section 5] if the Aubin property can be characterized by an exact form of a certain second-order optimality condition together with the constraint nondegeneracy. It is noted that our analysis for nonlinear semidefinite programming (and also for nonlinear second-order cone programming in [7]) relies on the explicit formulas of the coderivative for the underlying normal cone mapping. Currently, it is not clear to us how to extend these results to the generic non-polyhedral $C^{2}$-cone reducible constrained optimization problems. We leave this as our future research topic.

### v2

@@ 4 Characterizations of the Aubin property for NLSDP
It should be emphasized that, for (1.1) with ${\mathcal{K}}$ being an arbitrary $C^{2}$-cone reducible set, it is still unknown if the strong regularity of the KKT system (1.7) is equivalent to the constraint nondegeneracy (2.13) combined with a certain second-order optimality condition similar to (2.18). The currently known cases include the nonlinear programming [13], the NLSOCP [5], the NLSDP problem [45] and a composite matrix programming regarding matrix eigenvectors [9].
@@ 5 Conclusions
In this paper, we prove that at a locally optimal solution to the nonlinear semidefinite programming problem (1.8), the Aubin property of ${\mathbb{S}}_{\rm KKT}$ (1.6) is equivalent to the strong second-order sufficient condition plus the constraint nondegeneracy. This enables us to derive a series of equivalent characterizations of the Aubin property, which includes the strong regularity of the Karush-Kuhn-Tucker system (1.3). As a byproduct, for nonlinear semidefinite programming, this paper answers the open question posed in [11, Section 5] if the Aubin property can be characterized by an exact form of a certain second-order optimality condition together with the constraint nondegeneracy. It should be noted that our analysis for nonlinear semidefinite programming (and also for nonlinear second-order cone programming in [8]) relies on the explicit formulas of the coderivative for the underlying normal cone mapping. Currently, it is not clear to us how to extend these results to generic non-polyhedral $C^{2}$-cone reducible constrained optimization problems. We leave this as our future research topic.